### v4

@@ 6.2 Polychoric correlation matrix
Figure 5 visualizes the absolute average bias as well as confidence interval coverage at 95% nominal level (calculated over the 5,000 repetitions) of the robust estimator and the MLE for each pairwise polychoric correlation coefficient. As expected, when the model is correctly specified ($\varepsilon=0$), both estimators coincide with accurate estimates. However, in the presence of contamination ($\varepsilon>0$), the two estimators deviate. The MLE exhibits a notable bias for all correlation coefficients, which increases gradually with increasing contamination fraction. The magnitude of the bias tends to be larger for pairs with larger true correlation, such as 0.560.560.56 for pair (1,2)(1,2)( 1 , 2 ), than for pairs with weaker true correlation, such as 0.200.200.20 for pair (5,4)(5,4)( 5 , 4 ). In addition, coverage of the MLE drops quickly for many pairs and gradually for the remaining ones. Conversely, the robust estimator remains nearly unaffected for a broad range of contamination fractions for each correlation coefficient ($\varepsilon\leq 0.2$ or $\varepsilon\leq 0.3$ depending on the variable pair), with bias only somewhat increasing afterwards. Furthermore, its coverage remains close to 0.9 or higher even at the high contamination fraction of $\varepsilon=0.3$. This reflects excellent performance of the proposed estimator with respect to robustness to uninformative responses. Appendix D.2 contains additional evaluations. In essence, the robust estimator yields accurate standard errors, but its confidence intervals tend to be wider than those of the MLE in the presence of contamination.
@@ 6.3 Discussion and additional simulation experiments
The second experiment, described in Appendix E.3, focuses on correlation shifts in the contamination distribution $H$. Specifically, the contamination distribution is the same as the true model distribution except for a sign-flipped correlation coefficient $-\rho_{*}$. For moderate correlation $\rho_{*}$, the proposed estimator does not improve upon the MLE due to substantial overlap between the true model distribution and the contamination. However, the gain in robustness increases substantially for higher correlation coefficients $\rho_{*}$.
@@ 7.2 Results
To further investigate the presence of careless respondents who attenuate correlational estimates, we study in detail the adjective pair “not envious” and “envious”, which featured the largest discrepancy between the ML estimate and the robust estimate in Figure 6, with an absolute difference of 0.314. The results of the two estimators and, for completeness, the sample correlation, are summarized in Table 3. The ML estimate of 0.618-0.618- 0.618 and sample correlation estimate of 0.562-0.562- 0.562 for the (polychoric) correlation coefficient seem remarkably weak considering that the two adjectives in question are polar opposites. In contrast, the robust correlation estimate is given by 0.925-0.925- 0.925, which seems much more in line with what one would expect if all participants responded accurately and attentively (cf., Arias et al.,, 2020).
[…]
@@
As a robustness check on the role of tuning constant $c$, Figure 7 visualizes the point estimate $\widehat{\rho}_{N}$ of the polychoric correlation between the item pair “not envious” and “envious” for various values of $c$. The point estimate stays relatively constant for $c$ between 0 and 0.75, with $\widehat{\rho}_{N}$ ranging between 0.95-0.95- 0.95 and 0.92-0.92- 0.92. Just after $c=0.75$, $\widehat{\rho}_{N}$ abruptly jumps to about 0.85-0.85- 0.85, before it stabilizes again and slowly transitions to the ML estimate of 0.62-0.62- 0.62 (see Table 3) for $c\rightarrow\infty$. Since $\widehat{\rho}_{N}$ very slowly approaches the value of the ML estimate, we only visualize choices of $c$ up to 2 in Figure 7. The instability of the estimate around $c=0.75$ suggests that $c$ should be chosen below this value to avoid disproportionate influence of poorly fitting cells. For the broad range of $c\leq 0.75$, we obtain a robust finding of “not envious” and “envious” having a very strong negative correlation after accounting for likely careless responding.
@@
Overall, leveraging our robust estimator, we find evidence for the presence of careless respondents in the data of Arias et al., (2020). While they substantially affect the correlation estimate of the MLE, amounting to about 0.62-0.62- 0.62, which is much weaker than one would expect for polar opposite items, our robust estimator can withstand their influence with an estimate of about 0.93-0.93- 0.93 and also help identify them through unreasonably large PR values. On a final note, our findings from the empirical application align remarkably well with those of the simulation from Appendix E.3, therefore strenghening their validity. We provide a detailed discussion of this similarity in Appendix E.3.3.
@@ 8 Estimation under distributional misspecification
The simulation studies in Section 6 have been concerned with a situation in which the polychoric model is misspecified for a subset of a sample (partial misspecification). However, as outlined in Section 4.3, another misspecification framework of interest is that of distributional misspecification where the model is misspecified for the entire sample. Suppose that instead of a bivariate standard normal distribution, the latent variables $(\xi,\eta)$ jointly follow an unknown and unspecified distribution $G$. In this framework, the object of interest is the population correlation between $\xi$ and $\eta$ under distribution $G$, that is, $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, rather than a polychoric correlation coefficient. Estimators for such situations where the population distribution $G$ is nonnormal have been proposed by Lyhagen & Ornstein, (2023), Jin & Yang-Wallentin, (2017), and Roscino & Pollice, (2006).
@@ 8.2 Simulation study
To simulate ordinal variables that were generated by a nonnormal latent distribution $G$, we employ the VITA simulation method of Grønneberg & Foldnes, (2017). For a pre-specified value of the population correlation $\rho_{G}$, the VITA method models the latent random vector $(\xi,\eta)$ such that the individual variables $\xi$ and $\eta$ both possess standard normal marginal distributions with population correlation set equal to $\rho_{G}=\mathrm{Cor}_{G}\left[\xi,\ \eta\right]$, but are not jointly normally distributed. Instead, their joint distribution $G$ is equal to a pre-specified nonnormal copula distribution, such as the Clayton or Gumbel copula. Grønneberg & Foldnes, (2017) show that discretizing such VITA latent variables yields ordinal observations that could not have been generated by a standard bivariate normal distribution, thereby ensuring proper violation of the latent normality assumption.
@@
To investigate the robustness of our estimator to distributional misspecification, we use the VITA method to generate draws of the latent variables $(\xi,\eta)$ such that the latent variables are jointly distributed according to a Clayton copula $G$ with population correlation $\rho_{G}\in\{0.9,0.3\}$ (see Figure 8 for visualizations). Following the discretization process (1), we discretize both latent variables via discretization thresholds
@@
Figure 9 visualizes the bias of the robust estimator and the polychoric MLE under both Clayton copulas across the repetitions. [^15] In 27 of the 5,000 repetitions for the Clayton copula, the robust estimator experienced numerical instability and subsequently did not converge to a solution. The polychoric MLE failed to converge one time. Consequently, these unconverged estimates are omitted from Figure 9. We explain this numerical issue in detail in Appendix B. For correlation 0.9, the polychoric MLE exhibits a noteworthy bias, whereas the robust estimator remains accurate, albeit with a larger estimation variance as compared to partial misspecification (cf. Section 6). Conversely, for the weaker correlation 0.3, both estimators are fairly accurate with average biases of about 0.025-0.025- 0.025. Table 5 contains additional performance measures regarding inference. For $\rho_{G}=0.3$, the average standard error estimate of the robust method is accurate, but for $\rho_{G}=0.9$, it notably overestimates. We therefore also computed the median of the standard error estimates in the latter case: at 0.059, it still overestimates the approximate true standard error of 0.037, but far less so than the average. It turns out that there is a small number of simulated datasets with a large majority of empty cells in the contingency table, resulting in numerical instability of the standard error estimates and inflating their average. As discussed in Section 4.3, distributional misspecification is not covered by our partial misspecification framework, so it is not surprising that standard errors derived under partial misspecification are not always valid. Bootstrap inference may therefore be an attractive alternative. Nevertheless, unlike the polychoric MLE with zero coverage for $\rho_{G}=0.9$, the robust estimator maintains high coverage of over 90%.
@@ 9 Discussion and conclusion
However, our estimator depends on a user-specified choice of a tuning constant $c$, which governs a tradeoff between robustness and efficiency in case of misspecification. While simulation experiments suggest that the choice $c=0.6$ provides a good tradeoff and estimates do not change considerably for a broad range of finite choices of $c$, a detailed investigation on this tuning constant needs to be carried out in future work. As an alternative, one could consider other discrepancy functions that do not depend on tuning constants, like the ones discussed in Welz, (2024).
@@
The methodology proposed in this paper suggests a number of extensions. For instance, one could use a robustly estimated polychoric correlation matrix in structural equation models to robustify such models against misspecification. Similar robustification could by achieved in, for instance but not limited to, principal component analyses, multidimensional scaling, or clustering. In addition, the theory of Welz, (2024) may allow to pinpoint possible sources of model misspecification on the individual response level. That is, it may enable the derivation of statistically sound cutoff values for Pearson residuals in order to detect whether a given response can be fitted well by the polychoric correlation model. We leave these avenues to further research.
@@ Appendix B Remark on numerical optimization
[…]
@@ C.2 Simulation design
[…]
@@ C.3 Performance measures
[…]
@@ E.1 The role of distributional overlap
In general and beyond polychoric correlation, substantial overlap between the the model distribution and the contamination distribution in the contamination model of Huber, (1964), a particular identification issue may arise. For ease of exposition, even though similar arguments hold for estimating polychoric correlation, we explain this identification issue within the broader context of covariance matrix estimation. In this setting, said identification issue affects any affine equivariant robust estimator of the covariance matrix. Consider first a contamination fraction $\varepsilon=0.5$ and a contamination distribution $H$ so that we obtain two perfectly separated point clouds of exactly the same size. An affine equivariant robust estimator would have no way of knowing whether it should return the covariance parameters of the first group or the second group (cf. Section 6.17.1 in Maronna et al.,, 2018), which is the reason why the contamination fraction $\varepsilon$ is typically restricted to the interval [0,0.5)[0,0.5)[ 0 , 0.5 ) in the Huber contamination model. If the point clouds are no longer separated but overlap, fewer points from the groups are separated such that the identification issue already occurs for contamination fractions $\varepsilon$ of less than 0.5. This is because an affine equivariant covariance matrix estimator may successfully identify some points as outliers and downweight them, but it would have no way of identifying whether observations from the overlapping part come from the model distribution of interest or the contamination distribution, causing bias in the estimates. Riani et al., (2014) have formalized and studied this issue for various robust estimators in a regression setting.

### v5

@@ 6.2 Polychoric correlation matrix
Figure 5 visualizes the absolute average bias as well as confidence interval coverage at 95% nominal level (calculated over the 5,000 repetitions) of the robust estimator and the MLE for each pairwise polychoric correlation coefficient. As expected, when the model is correctly specified ($\varepsilon=0$), both estimators coincide with accurate estimates. However, in the presence of contamination ($\varepsilon>0$), the two estimators deviate. The MLE exhibits a notable bias for all correlation coefficients, which increases gradually with increasing contamination fraction. The magnitude of the bias tends to be larger for pairs with larger true correlation, such as $0.56$ for pair $(2,1)$, than for pairs with weaker true correlation, such as $0.20$ for pair $(5,4)$. In addition, coverage of the MLE drops quickly for many pairs and gradually for the remaining ones. Conversely, the robust estimator remains nearly unaffected for a broad range of contamination fractions for each correlation coefficient ($\varepsilon\leq 0.2$ or $\varepsilon\leq 0.3$ depending on the variable pair), with bias only somewhat increasing afterwards. Furthermore, its coverage remains close to 0.9 or higher even at the high contamination fraction of $\varepsilon=0.3$. This reflects excellent performance of the proposed estimator with respect to robustness to uninformative responses. Appendix D.2 contains additional evaluations. In essence, the robust estimator yields accurate standard errors, but its confidence intervals tend to be wider than those of the MLE in the presence of contamination.
We stress that polychoric correlation matrices need not be positive definite (e.g., Mair,, 2018, p. 22), although all estimated polychoric correlation matrices in this simulation turned out positive definite. If an estimated polychoric correlation matrix is not positive definite, one may opt to apply a smoothing procedure like in Yuan et al., (2011) or Bock et al., (1988).
@@ 6.3 Discussion and additional simulation experiments
The second experiment, described in Appendix E.3, focuses on correlation shifts in the contamination distribution $H$. Specifically, the contamination distribution is the same as the true model distribution except for a sign-flipped correlation coefficient $\rho_{*}$. For moderate correlation $\rho_{*}$, the proposed estimator does not improve upon the MLE due to substantial overlap between the true model distribution and the contamination. However, the gain in robustness increases substantially for higher correlation coefficients $\rho_{*}$. We expect the gain in robustness to increase for a higher number of response options and decrease for fewer response options. In the most extreme case of two dichotomous rating variables, no improvement can be expected.
@@ 7.2 Results
To further investigate the presence of careless respondents who attenuate correlational estimates, we study in detail the adjective pair “not envious” and “envious”, which featured the largest discrepancy between the ML estimate and the robust estimate in Figure 6, with an absolute difference of 0.314. The results of the two estimators and, for completeness, the sample correlation, are summarized in Table 3. The ML estimate of 0.618-0.618 and sample correlation estimate of 0.562-0.562 for the (polychoric) correlation coefficient seem remarkably weak considering that the two adjectives in question are polar opposites. In contrast, the robust correlation estimate is given by 0.925-0.925, which seems much more in line with what one would expect if all participants responded accurately and attentively (cf., Arias et al.,, 2020).
$X\YY$ 1 2 3 4 5
@@
As a robustness check on the role of tuning constant $c$, Figure 7 visualizes the point estimate $\widehat{\rho}_{N}$ of the polychoric correlation between the item pair “not envious” and “envious” for various values of $c$. The point estimate stays relatively constant for $c$ between 0 and 0.75, with $\widehat{\rho}_{N}$ ranging between 0.95-0.95 and 0.92-0.92. Just after $c=0.75$, $\widehat{\rho}_{N}$ abruptly jumps to about 0.85-0.85, before it stabilizes again and slowly transitions to the ML estimate of 0.62-0.62 (see Table 3) for $c\rightarrow\infty$. Since $\widehat{\rho}_{N}$ very slowly approaches the value of the ML estimate, we only visualize choices of $c$ up to 2 in Figure 7. The instability of the estimate around $c=0.75$ suggests that $c$ should be chosen below this value to avoid disproportionate influence of poorly fitting cells. For the broad range of $c\leq 0.75$, we obtain a robust finding of “not envious” and “envious” having a very strong negative correlation after accounting for likely careless responding.
@@
Overall, leveraging our robust estimator, we find evidence for the presence of careless respondents in the data of Arias et al., (2020). While they substantially affect the correlation estimate of the MLE, amounting to about 0.62-0.62, which is much weaker than one would expect for polar opposite items, our robust estimator can withstand their influence with an estimate of about 0.93-0.93 and also help identify them through unreasonably large PR values. On a final note, our findings from the empirical application align remarkably well with those of the simulation from Appendix E.3, therefore strenghening their validity. We provide a detailed discussion of this similarity in Appendix E.3.3.
@@ 8 Estimation under distributional misspecification
The simulation studies in Section 6 have been concerned with a situation in which the polychoric model is misspecified for a subset of a sample (partial misspecification). However, as outlined in Section 4.3, another misspecification framework of interest is that of distributional misspecification where the model is misspecified for the entire sample. Suppose that instead of a bivariate standard normal distribution, the latent variables $(\xi,\eta)$ jointly follow an unknown and unspecified distribution $G$. In this framework, the object of interest is the population correlation between $\xi$ and $\eta$ under distribution $G$, that is, $\rho_{G}=\mathbb{C}\mathrm{or}_{G}\left[\xi,\ \eta\right]$, rather than a polychoric correlation coefficient. Estimators for such situations where the population distribution $G$ is nonnormal have been proposed by Lyhagen & Ornstein, (2023), Jin & Yang-Wallentin, (2017), and Roscino & Pollice, (2006).
@@ 8.2 Simulation study
To simulate ordinal variables that were generated by a nonnormal latent distribution $G$, we employ the VITA simulation method of Grønneberg & Foldnes, (2017). For a pre-specified value of the population correlation $\rho_{G}$, the VITA method models the latent random vector $(\xi,\eta)$ such that the individual variables $\xi$ and $\eta$ both possess standard normal marginal distributions with population correlation set equal to $\rho_{G}=\mathbb{C}\mathrm{or}_{G}\left[\xi,\ \eta\right]$, but are not jointly normally distributed. Instead, their joint distribution $G$ is equal to a pre-specified nonnormal copula distribution, such as the Clayton or Gumbel copula. Grønneberg & Foldnes, (2017) show that discretizing such VITA latent variables yields ordinal observations that could not have been generated by a standard bivariate normal distribution, thereby ensuring proper violation of the latent normality assumption.
@@
To investigate the robustness of our estimator to distributional misspecification, we use the VITA implementation in package covsim (Grønneberg et al.,, 2022) to generate draws of the latent variables $(\xi,\eta)$ such that the latent variables are jointly distributed according to a Clayton copula $G$ with population correlation $\rho_{G}\in\{0.9,0.3\}$ (see Figure 8 for visualizations). Following the discretization process (1), we discretize both latent variables via discretization thresholds
$a_{1}={-1.5},$ $a_{2}={0},$ $a_{3}={0.5},$ $a_{4}={1},$  and
$b_{1}={-1},$ $b_{2}={1},$ $b_{3}={1.5},$ $b_{4}={2},$
@@
Figure 9 visualizes the bias of the robust estimator and the polychoric MLE under both Clayton copulas across the repetitions. [^15] In 32 of the 5,000 repetitions for the Clayton copula, the robust estimator experienced numerical instability and subsequently did not converge to a solution. The polychoric MLE failed to converge 56 times. Consequently, these unconverged estimates are omitted from Figure 9. We explain this numerical issue in detail in Appendix B. Note that due to the shape of the Clayton copula with high correlation, stability issues with the robust estimator are in our experience amplified if the true thresholds are not well distributed over the domain of the distribution. For correlation 0.9, the polychoric MLE exhibits a noteworthy bias, whereas the robust estimator remains accurate, albeit with a larger estimation variance as compared to most simulations configurations of partial misspecification (cf. Section 6). Conversely, for the weaker correlation 0.3, both estimators are fairly accurate with average biases of about 0.015-0.015. Table 5 contains additional performance measures regarding inference. For $\rho_{G}=0.3$, the average standard error estimate of the robust method is accurate, but for $\rho_{G}=0.9$, it notably overestimates. We therefore also computed the median of the standard error estimates in the latter case: at 0.017, it is fairly close to the true standard error of 0.014. It turns out that there is a small number of simulated datasets with a large majority of empty cells in the contingency table, resulting in numerical instability of the standard error estimates and inflating their average. As discussed in Section 4.3, distributional misspecification is not covered by our partial misspecification framework, so it is not surprising that standard errors derived under partial misspecification are not always valid. Bootstrap inference may therefore be an attractive alternative. Nevertheless, unlike the polychoric MLE with coverage of only about 13% for $\rho_{G}=0.9$, the robust estimator maintains high coverage of over 90%.
@@ 9 Discussion and conclusion
However, our estimator depends on a user-specified choice of a tuning constant $c$, which governs a tradeoff between robustness and efficiency in case of misspecification. While simulation experiments suggest that the choice $c=0.6$ provides a good tradeoff and estimates do not change considerably for a broad range of finite choices of $c$, a detailed investigation on this tuning constant needs to be carried out in future work. As an alternative, one could consider other discrepancy functions that do not depend on tuning constants, like the ones discussed in Welz, (2024). As practical guidelines, we recommend always comparing robust estimates to that of ML and running the robust estimator for various choices of $c$, like in Figure 7. Doing so not only helps assess the estimates’ stability, but also the severity of (partial) model misspecification. If the ML and robust estimates strongly differ, one may want to opt for choices of $c>0$ not too far from 0 to achieve larger robustness gains.
@@
The methodology proposed in this paper suggests a number of extensions. For instance, one could use a robustly estimated polychoric correlation matrix in structural equation models to robustify such models and their fit indices against misspecification. Similar robustification could by achieved in, for instance but not limited to, principal component analyses, multidimensional scaling, or clustering. In addition, the theory of Welz, (2024) may allow to pinpoint possible sources of model misspecification on the individual response level. That is, it may enable the derivation of statistically sound cutoff values for Pearson residuals in order to detect whether a given response can be fitted well by the polychoric correlation model. We leave these avenues to further research.
@@ Appendix B Remark on numerical optimization
In the simulation study on distributional misspecification from Section 8, the robust estimator did not converge to a solution in 32 out of 5,000 repetitions. This issue occurred 27 times for the Clayton copula with correlation 0.9, and five times with correlation 0.3. The polychoric MLE failed to converge 56 times, all at the Clayton copula with correlation 0.9. For the robust estimator, this situation tends to occur in the rare event that single rows or columns in a contingency table contain only one non-zero cell. It turns out that for such data, the estimator may attempt to effectively eliminate a threshold by either pushing the outermost finite thresholds to $\pm\infty$ or pulling adjacent thresholds to be as close to each other as numerically feasible, thereby practically merging these thresholds into one. Such behavior often leads to degeneracy of the final Nelder-Mead simplex (if used). Thus, package robcat identifies numerical instability either by degeneracy of Nelder-Mead simplexes and/or adjacent thresholds being unreasonably far away from each other, and subsequently throws a warning. We decided to set the cutoff for two adjacent thresholds being unreasonably far away to a minimum distance of 3.92. Under the polychoric model, this distance can cover as much as 95% of all probability mass of the corresponding standard normal marginal distribution. We believe that having adjacent threshold values separated by this much probability mass may be an indication of poor model fit and/or severe measurement issues. Hence, being able to detect such numerical instability may in fact be viewed as a useful feature of our robust estimator: it alerts the user to potential broader issues where the use of the model may not be recommended for the data at hand.
@@ C.2 Simulation design
To simulate contamination, we let a fraction $\varepsilon\in\{0,0.1,0.2\}$ of the data be generated by a contamination distribution $H$, which here is a bivariate normal distribution with population mean $(2.5,-2.5)^{\top}$, variances $(0.25,0.25)^{\top}$, and zero correlation between the two latent variables. We then discretize the latent realizations from this contamination distribution according to the thresholds $(a_{*,j},b_{*,j},j=1,\dots,4)$. We sample $N=1,000$ ordinal observations $(X_{i},Y_{i})_{i=1}^{N}$ from this data generating process. We then use this sample to estimate the polychoric correlation between the two ordinal variables with our proposed robust estimator at various choices of tuning parameter $c$. Specifically, the set of considered tuning parameters $c$ is given by the granular grid $\{0,0.1,0.2,\dots,14.8,14.9,15,+\infty\}$, where $c=+\infty$ is understood as the maximum likelihood estimator (MLE). We repeat this procedure for $T=5,000$ simulation runs.
@@ C.3 Performance measures
Comparing $\widehat{\textnormal{SE}}^{\textnormal{ave}}\left(\widehat{\rho}_{N}\right)$ with $\textnormal{SE}^{\textnormal{approx}}\left(\widehat{\rho}_{N}\right)$ can be seen as a sanity check for the correctness of the estimator’s limit theory in Theorem 1. Hence, if the limit theory is correct, we expect for a sufficiently large sample size that $\widehat{\textnormal{SE}}^{\textnormal{ave}}\left(\widehat{\rho}_{N}\right)$ will be close to $\textnormal{SE}^{\textnormal{approx}}\left(\widehat{\rho}_{N}\right)$. However, recall from Theorem 1 that the estimator is not asymptotically Gaussian for the choice $c=0$ when contamination is absent $(\varepsilon=0)$. Thus, in this case ($c=\varepsilon=0$), we expect $\widehat{\textnormal{SE}}^{\textnormal{ave}}\left(\widehat{\rho}_{N}\right)$ to differ from $\textnormal{SE}^{\textnormal{approx}}\left(\widehat{\rho}_{N}\right)$ because the former is derived under the then-invalid asymptotic normality of estimator $\widehat{\bm{\theta}}_{N}$.
@@ E.1 The role of distributional overlap
In general and beyond polychoric correlation, substantial overlap between the the model distribution and the contamination distribution in the contamination model of Huber, (1964), a particular identification issue may arise. For ease of exposition, even though similar arguments hold for estimating polychoric correlation, we explain this identification issue within the broader context of covariance matrix estimation. In this setting, said identification issue affects any affine equivariant robust estimator of the covariance matrix. Consider first a contamination fraction $\varepsilon=0.5$ and a contamination distribution $H$ so that we obtain two perfectly separated point clouds of exactly the same size. An affine equivariant robust estimator would have no way of knowing whether it should return the covariance parameters of the first group or the second group (cf. Section 6.17.1 in Maronna et al.,, 2018), which is the reason why the contamination fraction $\varepsilon$ is typically restricted to the interval $[0,0.5)$ in the Huber contamination model. If the point clouds are no longer separated but overlap, fewer points from the groups are separated such that the identification issue already occurs for contamination fractions $\varepsilon$ of less than 0.5. This is because an affine equivariant covariance matrix estimator may successfully identify some points as outliers and downweight them, but it would have no way of identifying whether observations from the overlapping part come from the model distribution of interest or the contamination distribution, causing bias in the estimates. Riani et al., (2014) have formalized and studied this issue for various robust estimators in a regression setting.